# Simulating methylamine using symmetry adapted qubit-excitation-based variational quantum eigensolver

### v2

@@ IV VQE simulations
The use of more complex bases that allow achieving better description of valence orbitals or account for polarization effects and electron system correlations in molecules, such as 6-31g, 6-31g*, or cc-pVDZ, is associated with similar difficulties. All of them imply that significantly more qubits are required compared to the case of the minimal basis. For instance, simulating the H4 molecule using the cc-pVDZ basis would require 40 qubits Guo et al. (2023). For simulating the H2O molecule in the 6-31g basis, 50 qubits are needed. For methylamine and formic acid molecules the number of qubits exceed 60 in the case of complex bases.
Recent studies employing density matrix embedding methods in conjunction with the VQE algorithm demonstrate the simulation of relatively large systems with small computational resources. For instance, the simulation of the C18 molecule (a model system consisting of 144 spin orbitals and 72 electrons) has been conducted on a 16-qubit system et al. (2021). The simulation of the torsional barrier of ethane and the quantification of protein-ligand interactions has been demonstrated Shang et al. (2023), and the comparison of the use of High-Performance Computing and Matrix Product States for the ansatzes has been presented. Despite the size of the electronic systems in such studies, they often employ the HOMO-LUMO approach to reduce the active space of available orbitals. This leads to a more compact system representation in terms of qubits but comes at the cost of additional error.
We demonstrate the results of simulations for the Methylamine molecule with a restricted number of spin orbitals. We take an active space of 6 electrons and 6 spin orbitals from the full 15 spin-orbital space. First, we will show the energy difference for such a choice. In Fig. 3, one can observe different configurations of the methylamine active space and their corresponding energies. This choice of active space reduces the number of qubits needed for the molecular system. However, when the number of spin orbitals decreases, the calculated energy becomes closer to the Hartree-Fock results. The chemical accuracy is achieved only when the core 1s orbitals are frozen.
For the (6e, 6o) active space configuration of methylamine we use the UCCSDQ and UCCSDQs ansatzes in VQE calculation and the same minimal basis STO-3G. As it shown in Fig 4 the energy converged to the FCI value for both ansatzes. This result shows that for 12-qubit methylamine active space VQE in combination with aforementioned combination of the optimization methods could achieve the chemical accuracy within small number of iterations. Other types of ansatzes were not tested due to worser computational cost.
The VQE calculation for the formic acid molecule in a reduced active space, similar to the one performed for the methylamine, did not show good convergence to the expected ground state energy. The reason might lie in the stronger electronic correlations between carbon and oxygen atoms in the formic acid, which could not be adequately modeled by circuits based on qubit excitations. It is possible that effective fermionic excitation circuits, as proposed in Ref. Yordanov et al. (2020), could achieve better performance, however at the cost of an increased two-qubit gate count.
@@ V Conclusions
Quantum chemistry is one of the most interesting avenues for applications of quantum computing devices. In this work, we have presented the results of the computational resources analysis for certain molecules, such as formic acid and methylamine. We have considered the VQE approach and its various extensions as the primary method for quantum chemical simulation.
In order to illustrate the VQE approach, we have demonstrated energy spectra calculations for small molecules, such as LiH and BeH2222, and for methylamine with resticted active space of spin orbitals, all in the minimal basis. Our analysis has shown that the combination of qubit excitations with filtration by irreducible representations (UCCSDQs) could drastically decrease the ansatz depth and the number of two-qubit gates, while still recovering the major part of the correlation energy. Combining these methods with the qubit tapering procedure could lead to further improvements; however, it requires a more careful analysis of the connection between Z2 symmetries and the representation of the excitation operator circuit, which is beyond the scope of this study. The k-UpCCGSD ansatz could result in smaller improvements, but it can be easily combined with qubit tapering. Its combination with irreducible representation filtering showed poor convergence. The ability of these ansatzes to avoid barren plateaus Wang et al. (2021) was not investigated in our study. A possible scenario is that, as the number of qubits increases, the process of classical optimization could become stuck in local minima, in the case of the k-UpCCGSD ansatz it could potentially be fixed by increasing $k$, i.e., by increasing the number of variational parameters.
The quality of two-qubit operations in quantum computing devices is a crucial bottleneck in the real-world application of quantum chemical simulations, e.g. with the use of the using naive UCCSD ansatzes. We have considered possible ways for overcoming this challenges by modifying the algorithmic approach. For example, the combination of k-UpUCCGSD with qubit excitations can significantly decrease the circuit depth and the number of two-qubit gates. The addition of irreducible representation filtering (k-UpGSDQs) could further almost halve the costs. However, the ability to achieve accurate results with these combinations for methylamine and formic acid in the full active space of spin orbitals remains a subject for further study.
As we expect, further improvements in ansatz design, gate fidelity, and error mitigation may greatly lessen these restrictions and open the door for real-world uses of quantum algorithms like VQE in quantum chemistry.

### v3

@@ IV Results
The VQE calculation for the formic acid molecule in a reduced active space, similar to the one performed for methylamine, does not show good convergence to the expected ground state energy. The reason might lie in the stronger electronic correlations between carbon and oxygen atoms in the formic acid molecule, which could not be adequately modeled by circuits based on qubit excitations. It is possible that effective fermionic excitation circuits, as proposed in Ref. Yordanov et al. (2020), could achieve better performance, however at the cost of an increased two-qubit gate count.
@@ V Conclusions
Quantum chemistry is one of the most promising avenues for applications of quantum computing devices, with the VQE and its various extensions currently serving as the primary methods for quantum chemical simulations. In this work, we present the results of a computational resource analysis for organic molecules, such as formic acid and methylamine, within the VQE framework. Additionally, we investigate several optimization strategies to achieve better quantum resource scaling. To the best of our knowledge, no prior studies have systematically explored the combined effect of these optimization strategies on VQE performance.
To demonstrate their adequacy, we calculate the energy spectra of small molecules, such as LiH and BeH2, as well as methylamine with a restricted active space of spin orbitals, all in the minimal basis. Our analysis show that the combination of qubit excitations with filtration by irreducible representations (UCCSDQs) could drastically decrease the ansatz depth and the number of two-qubit gates, while still recovering the major part of the correlation energy.
We estimate the quantum resources required for full active-space configuration simulations of the methylamine and formic acid molecules using different types of ansatzes and present the results in Table 2 for comparison. We highlight the UCCSDQs results as most promising optimization strategy in terms of the circuit depth and algorithm convergence. Combining UCCSDQs with the qubit tapering procedure could lead to further improvements; however, it requires a more careful analysis of the connection between $\mathbb{Z}_{2}$ symmetries and the representation of the excitation operator circuits, which is beyond the scope of this study. The k-UpGSD ansatz may result in smaller improvements, but at the same time, it can be easily combined with qubit tapering. A possible challenge for this ansatz is that, as the number of qubits increases, the process of classical optimization may become stuck in local minima. In such cases, this issue could potentially be mitigated by increasing $k$, i.e., by increasing the number of variational parameters. For example, in the BeH2 calculation (see Fig. 2), we used $k=4$ to achieve convergence. This, however, comes at the cost of increased circuit depth, which amplifies the impact of two-qubit gate errors.
The quality of two-qubit operations in quantum computing devices is a crucial bottleneck for the real-world application of quantum chemical simulations. As a result, ansatzes that require a large number of such operations become impractical, necessitating the search for optimization strategies. We have considered possible ways for overcoming these challenges by using various combinations of previously developed optimization methods. However, challenges remain, particularly for larger molecules and more complex systems. We expect that further improvements in ansatz design, gate fidelity, and error mitigation may greatly lessen these restrictions and open the door for practical applications of quantum algorithms such as VQE in quantum chemistry.
Data availability. The data supporting the findings of this study are available from the corresponding authors upon reasonable request.
@@
Competing interests. Owing to the employments and consulting activities of K.M.M. and A.K.F., they have financial interests in the commercial applications of quantum computing.